OpenStudy (anonymous):

hat are the solutions of the system? y=x^2-2x-1 y=-2x

OpenStudy (dumbcow):

-2x = x^2 -2x -1

OpenStudy (anonymous):

Put the two x equations equal two each other, so -2x = x^2-2x-1. Then solve as a quadtratic so x^2-1=0 x = +1 and -1 Put that back into one of the original to get y=-2(1) and y=-2(-1) So y = -2 or 2 Final answer is (1,-2) and (-1,2)